# Quantize What Counts: Bit Allocation Insights Informed by Spectral Gaps in Keys and Values

### v2

@@ 3.2 Key-Driven Quantization
Consider an additive-residual Transformer block (Layer Norm omitted for clarity [37]),
$h_{l+1}=h_{l}+W_{l}h_{l},$
and let $\Delta_{l}$ denote the quantization perturbation applied to $W_{l}$. A single block incurs
@@
After $L$ layers the worst-case deviation satisfies
@@
From  Key-Value Norm Disparity:
$\mathbb{E}\!\bigl{[}\|W^{K}\|_{F}^{2}\bigr{]}>\mathbb{E}\!\bigl{[}\|W^{V}\|_{F}^{2}\bigr{]}.$ (i)
Hence any distortion injected on the key path is amplified more strongly throughout the layers.
Let $X\in\mathbb{R}^{\mathrm{seq}\times d_{\mathrm{model}}}$ be the hidden-state matrix at a given layer. The key and value caches are produced by the same input but different projections:
@@
Because $X$ is common to both paths, multiplying (i) on the left by $X$ and using the sub-multiplicativity of the Frobenius norm, $\|AB\|_{F}\leq\|A\|_{F}\,\|B\|_{2}$, yields
$\mathbb{E}\!\bigl{[}\|K\|_{F}^{2}\bigr{]}>\mathbb{E}\!\bigl{[}\|V\|_{F}^{2}\bigr{]},$ (ii)
i.e. the norm gap persists in the caches themselves.
Quantization error vs. Frobenius norm.
@@
For a matrix $M\in\mathbb{R}^{m\times n}$ the squared Frobenius norm $\|M\|_{F}^{2}$ equals the total energy $\sum_{i,j}M_{ij}^{2}$, or equivalently the sum of eigenvalues of $M^{\top}M$ [^1] Because $\operatorname{tr}(AB)=\operatorname{tr}(BA)$, one has $\operatorname{tr}\!\bigl{(}M^{\top}M\bigr{)}=\operatorname{tr}\!\bigl{(}MM^{\top}\bigr{)}$. The non-zero eigenvalues of $M^{\top}M$ and $MM^{\top}$ are identical and equal to the squared singular values $\sigma_{i}^{2}$ of $M$; hence $\|M\|_{F}^{2}=\sum_{i}\sigma_{i}^{2}=\operatorname{tr}\!\bigl{(}M^{\top}M\bigr{)}$.. When $M$ is quantized to $\tilde{M}$ with $b$ bits, uniform-scalar quantizers yield:
[…]
where [^2] Big-$\Theta$ notation: $f(n)=\Theta(g(n))$ iff $\exists\,c_{1},c_{2}>0,\;n_{0}$ such that $c_{1}g(n)\leq f(n)\leq c_{2}g(n)$ for all $n\geq n_{0}$. the hidden constants depend only on the quantizer variant. Since Key and Value caches share the same shape, minimizing mean square error reduces to minimizing $\|\,\cdot\,\|_{F}^{2}$. When $\|K\|_{F}\gg\|V\|_{F}$ and equal bit-widths are used,
$\mathbb{E}\!\bigl{[}\|K-\tilde{K}\|_{F}^{2}\bigr{]}\;\gg\;\mathbb{E}\!\bigl{[}\|V-\tilde{V}\|_{F}^{2}\bigr{]},$
making the key cache the dominant source of degradation, formally:
Theorem 2 (Key-Driven Quantization).
@@
Let $(b_{K},b_{V})$ denote the bit allocations assigned to the key and value caches, respectively, under a symmetric uniform-scalar quantizer. For any pair $(b_{K},b_{V})$ and $b_{K}>b_{V}$, the expected inference accuracy is strictly higher than for the swapped allocation $(b_{V},b_{K})$, provided that $\mathbb{E}\!\bigl{[}\|K\|_{F}^{2}\bigr{]}>\mathbb{E}\!\bigl{[}\|V\|_{F}^{2}\bigr{]}$.
Figure 3: Singular value distribution of key and value activations for Llama 3.3-70B on C4. The x-axis represents the singular value indices, ordered from the 5th largest to the smallest, while the y-axis denotes the magnitude of the singular values. The shaded region indicates the minimum-to-maximum range across attention heads within each layer, and the solid lines represent the mean singular value magnitude at each index. This zoomed-in view reveals that, beyond the largest singular value (spectral norm), key matrices exhibit consistently higher singular values even in the lower ranges, underscoring their greater representational significance compared to value matrices. For the complete range, see Figure 5 in Appendix E.
@@
Appendix 3 and  4 establishes a norm-dependent upper bound on the quantization error of an arbitrary matrix $M$, and Appendix C.6 applies this bound to show that key caches incur proportionally larger errors than value caches.
To empirically validate the theoretical predictions, we examined the singular value spectra of the K anc V caches across a broad set of models, up to 70B, by extracting every attention head in every layer (Section 4.1). Figure 3 displays the singular values from the fifth largest onward, revealing a systematic separation between keys and values; the complete spectra in Appendix E (Figure 5) show the same trend across all modes. Aggregate norms reinforce this observation: Figures 7 and 7 plot the layer-wise Frobenius and spectral norms of the $K\!V$ cache in Llama-3.3-70B, consistently indicating larger magnitudes for keys than for values.
4 Experimental Results
4.1 Experiment Setup
We conduct two experiments: (i) a quantization-error analysis and (ii) a downstream evaluation. For the quantization-error analysis, we use PyTorch’s weight-packing quantization to measure reconstruction error. For the downstream evaluation, we employ hardware-aligned quantization via Hugging Face’s Optimum Quanto library as back-end (token-wise, 2/4-bit quantization) on Llama-3.2-1B, Llama-3.1-8B [38], Phi-4-14B [39] and DeepSeek-R1-Qwen-14B, and the Half-Quadratic Quantization (HQQ) [40] back-end (channel-wise, 1–8-bit quantization) on Llama-3.1-8B, Llama-3.2-1B, Llama-3.2-3B, and Qwen3-0.6B/1.7B/4B/8B [41]. The quantization-error sweep additionally includes Phi-3-Medium-128K [42], Llama-3.1-Nemotron-70B [43], Mistral-0.3-7B [8], Qwen-2.5-14B and Llama-3.3-70B. See Appendix G for full details on compute resources, software tools, and reproducibility.
Table 1: Quantization errors for KV cache under various configurations. The quantization error is expressed as mean ± std. We use 2-bit, 3-bit, and 4-bit quantization for both $K_{i}$ and $V_{i}$ across three datasets.
@@ 4.1 Experiment Setup
$K_{2}$ $V_{2}$ $K_{3}$ $V_{3}$ $K_{4}$ $V_{4}$
@@
4.2 Quantization Error
In several studies (Table 3 of KIVI[29], Table 8 of GEAR[44], and Table 4 of RotateKV[45]), quantized KV cache representations have been shown to outperform their BF16 full-precision counterparts, an outcome that may seem counterintuitive. Taking BF16 as the gold standard, we attribute this phenomenon to the implicit regularization effect introduced by quantization: by discretizing activation values, quantization injects a small, structured noise into the model’s intermediate representations. Much like dropout or weight decay, this noise can smooth overly sharp decision boundaries and mitigate over-fitting, thereby improving generalization on held-out tasks. To isolate and measure this effect, our first experiment focuses on characterizing the quantization error in the KV cache, without grouping tokens, or using residual tokens in each quantization turn.
In addition to per layer and per head comparison, we average MSE across all heads, layers, and tokens in a batch. This approach yields a global view of how severely each bit-width compresses $\mathbf{K}$ vs. $\mathbf{V}$.
The evaluation uses ten random sequences from C4 [46], MMLU [47], and GSM8K [48], padded to the longest length and generated autoregressively for up to 1 000 tokens. Table 1 presents 2-, 3-, and 4-bit errors for seven representative models (Llama3, Phi4 [39], Mistral, Qwen, DeepSeek-Distilled, and others); the complete results appear in Appendix F. Figure 8 plots the full error curve for Llama-3.3-70B on C4, while Figure 10 shows the per-layer breakdown for Llama-3.1-8B.
4.3 Evaluation
Experiments in Section 4.2 established that K cache incur larger MSE than V cache at every bit-width, suggesting that $\mathbf{K}$ contains proportionally richer information. However, ultimately model quality is measured by downstream metrics such as accuracy rather than reconstruction error. Our second experiments therefore measures how the observed error asymmetry translates into task performance under realistic, deployment-oriented constraints.
KV cache quantization methods rarely de-quantize the entire cache after every generated token because the associated memory traffic defeats the purpose of quantization. Instead, they maintain a residual buffer that holds the most recent tokens in full-precision and flushes it en bloc once full. We replicate this strategy by (i) reserving a fixed-size residual cache of 64 tokens and (ii) grouping activations into 128-element blocks—settings adopted by popular open-source implementations such as KIVI, Flash-Decoding [49] and Marlin [50]. Two hardware-aligned quantization kernels were selected to span the design space:
  • Quanto: token-wise (per-row) quantization applied to the key and value caches, supporting $\{2,4\}$-bit precision.
  • HQQ: channel-wise (per-column) quantization applied to the key and value caches, supporting $\{1,2,4,6,8\}$-bit precision.
For downstream evaluation, we employ three generative datasets that require the model to produce extended outputs rather than select among fixed choices: the GSM8K, the CoQA [51] conversational question answering dataset, and the EQ-Bench [52] long-form structured QA suite.
Results.
Tables 2 and 3 corroborate the norm-based hypothesis by showing that downstream accuracy is far more sensitive to the precision assigned to the key cache than to the value cache.
  • Quanto backend. Across four models and two prompt lengths, allocating $4$ bits to keys and $2$ bits to values ($\mathbf{K}_{4}\mathbf{V}_{2}$) recovers on average $94\%$ of the full-precision score, whereas the inverted scheme ($\mathbf{K}_{2}\mathbf{V}_{4}$) loses up to $30$ percentage points (pp). The margin widens with model capacity: in 1-shot GSM8K, $\mathbf{K}_{4}\mathbf{V}_{2}$ outperforms $\mathbf{K}_{2}\mathbf{V}_{4}$ by $+0.30$ pp on Llama-3.2-1B and by $+0.16$ pp on Phi-4-14B. Notably, $\mathbf{K}_{4}\mathbf{V}_{2}$ attains almost the same accuracy as the symmetric $4$-bit baseline.
  • HQQ backend. The broader sweep over $\{1,2,4,6,8\}$ bits and three datasets paints the same picture. On GSM8K, granting even a single extra bit to keys yields double-digit gains; e.g., Qwen3-8B improves from $0.08$ to $0.54$ accuracy ($+46$ pp) when the extra bit is moved from values to keys. Gaps persist on EQ-Bench and CoQA, where key-centric quantization delivers up to $+62$ pp (Qwen3-0.6B, EQ-Parseable, $1$-bit) and never underperforms the converse setting. Advantages taper but do not vanish at higher precisions: at $6$ bits the key-first policy still adds $7\!-\!12$ pp for 8–14B models.
Table 2: KV cache quantization on GSM8K 1-shot and 8-shot performance employing Optimum Quanto backend. $\mathbf{K}_{i}\mathbf{V}_{j}$ denotes $i$-bit quantization for the K cache and $j$-bit for the V cache using Quanto backend. The $\mathbf{K}_{\mathbf{(4bit)}}\mathbf{V}_{\mathbf{(2bit)}}$ column, highlighted below, shows the optimal bit allocation setting.
1-shot 8-shot
Model $\mathbf{K}_{\mathbf{2}}\mathbf{V}_{\mathbf{2}}$ $\mathbf{K}_{\mathbf{2}}\mathbf{V}_{\mathbf{4}}$ $\mathbf{K}_{\mathbf{4}}\mathbf{V}_{\mathbf{2}}$ $\mathbf{K}_{\mathbf{4}}\mathbf{V}_{\mathbf{4}}$ $\mathbf{K}_{\mathbf{2}}\mathbf{V}_{\mathbf{2}}$ $\mathbf{K}_{\mathbf{2}}\mathbf{V}_{\mathbf{4}}$ $\mathbf{K}_{\mathbf{4}}\mathbf{V}_{\mathbf{2}}$ $\mathbf{K}_{\mathbf{4}}\mathbf{V}_{\mathbf{4}}$
Llama3.2-1B-it 0.033 0.035 0.338 0.357 0.031 0.031 0.289 0.369
Llama3.1-8B-it 0.511 0.547 0.752 0.754 0.408 0.441 0.770 0.782
Phi4-14B 0.759 0.783 0.913 0.923 0.771 0.815 0.927 0.931
DeepSeekR1Q-14B 0.772 0.775 0.865 0.867 0.763 0.792 0.876 0.875
22footnotetext: Although Chain-of-Thought prompting is generally discouraged for reasoning models and can even impair their performance, we deliberately apply 1-shot and 5-shot CoT prompts here to examine the effect of longer input prompts on KV cache quantization.
Table 3: KV cache quantization comparison across GSM8K, EQ-Parseable, and CoQA datasets. The results are obtained on the HQQ quantization backend. Each cell shows accuracies for $\mathbf{K}_{i}\mathbf{V}_{x}$ and $\mathbf{K}_{x}\mathbf{V}_{i}$, where $x$ denotes the mean accuracy across quantization bit-widths $\{1,2,4,6,8\}$, with the percentage-point gap annotated above the comparison sign. See Appendix H for all settings.

### v3

@@ 3.2 Key-Prioritized Quantization: Key-Favored Allocation Minimizes Quantization Error
Consider an additive-residual Transformer block (layer normalization omitted for clarity Elhage et al. (2021)):
$h_{l+1}=h_{l}+W_{l}h_{l}.$
Quantizing $W_{l}$ to $\tilde{W}_{l}=W_{l}+\Delta_{l}$ introduces an error bounded by
@@
After $L$ layers, the worst-case deviation accumulates multiplicatively:
@@
Because the expected norms satisfy
$\mathbb{E}\bigl[\|W^{K}\|_{F}^{2}\bigr]>\mathbb{E}\bigl[\|W^{V}\|_{F}^{2}\bigr],$
any quantization noise injected along the key path is amplified more strongly through the network.
Let $X\in\mathbb{R}^{\text{seq}\times d_{\text{model}}}$ be the hidden-state matrix at a given layer. The same input generates both caches via
@@
Multiplying the previous inequality by $X$ and applying sub-multiplicativity of the Frobenius norm yields
$\mathbb{E}\bigl[\|K\|_{F}^{2}\bigr]>\mathbb{E}\bigl[\|V\|_{F}^{2}\bigr],$
i.e., the norm gap persists in the caches.
Quantization Error and Norm Magnitude.
@@
For a matrix $M\in\mathbb{R}^{m\times n}$, the squared Frobenius norm equals the total signal energy and the sum of squared singular values. When quantized with $b$-bit uniform scalar quantization, the expected mean-square error satisfies
$\mathbb{E}\bigl[\|M-\tilde{M}\|_{F}^{2}\bigr]=\Theta\bigl(\|M\|_{F}^{2}\,2^{-2b}\bigr),$
with constants depending only on the quantizer. Since key and value caches have identical shape, minimizing quantization error reduces to allocating bits in proportion to their energy. When $\|K\|_{F}\gg\|V\|_{F}$ and equal bit-widths are used, key-cache error dominates:
$\mathbb{E}\bigl[\|K-\tilde{K}\|_{F}^{2}\bigr]\gg\mathbb{E}\bigl[\|V-\tilde{V}\|_{F}^{2}\bigr].$
This asymmetry in quantization error directly motivates an asymmetric bit allocation strategy, formalized as the following theorem:
@@
Theorem 2 (Key-Prioritized Quantization) Let $(b_{K},b_{V})$ denote the bit allocations for key and value caches under a uniform scalar quantizer. For any pair with $b_{K}>b_{V}$, the expected inference accuracy is strictly higher than for the swapped allocation $(b_{V},b_{K})$, provided that $\mathbb{E}[\|K\|_{F}^{2}]>\mathbb{E}[\|V\|_{F}^{2}]$.
Figure 3 provides empirical evidence for this effect. For Llama 3.3-70B on C4, the singular value spectra of the key caches consistently exceed those of the value caches beyond the top singular mode, indicating higher representational significance throughout the spectrum. Appendix C.3 extends this analysis to the full singular value range, where Figure 7 shows layer-wise Frobenius and spectral norms, all consistently revealing larger magnitudes for keys than for values.
@@
The theoretical underpinnings of this phenomenon are established in Theorems B.2 of Appendix B.2 and B.3 of Appendix B.3, which derive a norm-dependent upper bound on the quantization error of an arbitrary matrix M. Appendix B.4 then applies this bound to the key and value caches, showing that the larger norms of the key caches translate into proportionally higher quantization errors under equal bit allocations.
Figure 3: Singular value spectra of key and value activations in Llama 3.3-70B on C4 benchmark dataset. The x-axis shows singular value indices, ordered from the 5th largest onward for cleaner illustration, and the y-axis shows their magnitudes. Shaded regions mark the minimum-maximum range across attention heads within each layer, while dashed lines indicate the mean at each index. Beyond the top singular value (i.e., the spectral norm), key activations consistently exhibit larger singular values than value activations across the spectrum, highlighting their greater representational capacity. Full spectra are provided in Figure 6 of Appendix C.3.
4 Results
4.1 Experimental Setup
Three evaluations are conducted: (i) a quantization-error analysis, which measures reconstruction error across different bit-widths; (ii) a downstream task accuracy evaluation, which assesses how mixed-precision KV cache quantization affects model accuracy on practical benchmarks; and (iii) an integration case study with rotation-based outlier distribution methods, which investigates the effect of combining mixed precisions with rotation strategies. Full details on compute resources and software configurations are supplied in Appendix D.
Table 1: Quantization error (lower is better) of $K$ and $V$ caches under matched bit-widths. Each cell reports the mean $\pm$ standard deviation of the reconstruction MSE between the dequantized cache and its BF16 reference, averaged over layers, heads, and tokens (10 sequences per dataset up to 1,000 tokens each). Both caches are quantized to the same precision, i.e., $\mathbf{K}_{2}\mathbf{V}_{2}$, $\mathbf{K}_{3}\mathbf{V}_{3}$, and $\mathbf{K}_{4}\mathbf{V}_{4}$, to isolate their intrinsic sensitivity at equal bit budgets. Across model families (Llama, Phi, Mistral, Qwen, DeepSeek) and datasets (C4, MMLU, GSM8K), $\mathbf{K}$ consistently exhibits larger reconstruction error than $\mathbf{V}$ at the same bit-width, and the error decreases monotonically with increasing precision, indicating that keys are the dominant source of quantization distortion.
@@ 4.1 Experimental Setup
$\mathbf{K}_{\mathbf{2}}$ $\mathbf{V}_{\mathbf{2}}$ $\mathbf{K}_{\mathbf{3}}$ $\mathbf{V}_{\mathbf{3}}$ $\mathbf{K}_{\mathbf{4}}$ $\mathbf{V}_{\mathbf{4}}$
@@
Quantization-Error Evaluation. PyTorch’s weight-packing quantization is applied without residual buffers or activation grouping to isolate quantization effects. Ten random sequences are sampled from C4 Dodge et al. (2021), MMLU Hendrycks et al. (2021), and GSM8K Cobbe et al. (2021), padded to the longest length, and generated autoregressively for up to 1,000 tokens. Both per-layer and per-head reconstruction errors are computed, along with global averages across all heads, layers, and tokens, to offer a comprehensive view of bit-width sensitivity. This experiment spans seven models, including Llama-3.2-1B, Llama-3.1-8B Grattafiori et al. (2024), Phi-4-14B Abdin et al. (2024b), Mistral-0.3-7B Mistral AI (2024), Qwen-2.5-14B, Llama-3.3-70B, DeepSeek-R1-Qwen-14B, Phi-3-Medium-128K Abdin et al. (2024a), and Llama-3.1-Nemotron-70B Nvidia et al. (2024).
Downstream Task Accuracy. Two representative quantization backends are employed. Optimum Quanto applies token-wise (per-row) quantization with mixed 2/4-bit precision on Llama-3.2-1B, Llama-3.1-8B, Phi-4-14B, and DeepSeek-R1-Qwen-14B. HQQ Badri and Shaji (2023) applies channel-wise (per-column) quantization with bit-widths $\{1,2,4,6,8\}$ on Llama-3.1-8B, Llama-3.2-1B, Llama-3.2-3B, and Qwen3-0.6B/1.7B/4B/8B Qwen Team (2025). A 64-token residual buffer (which stores the most recent tokens in full precision before being periodically flushed) and 128-element activation grouping (which quantizes activations in fixed-size blocks to reduce overhead) are adopted to mirror practical decoding configurations used in KIVI Liu et al. (2024), Flash-Decoding Dao et al. (2023), and Marlin Frantar et al. (2025), ensuring that the evaluation reflects realistic deployment rather than idealized settings. Accuracy is measured on three generative benchmarks: GSM8K Cobbe et al. (2021), CoQA Reddy et al. (2019), and EQ-Bench Paech (2024), which collectively probe mathematical reasoning, conversational QA, and structured long-form generation.
Integration with Rotation-Based Methods. QuaRot Ashkboos et al. (2024) is selected for this case study. It applies structured randomized Hadamard rotations to activations before quantization, effectively dispersing outliers and improving the uniformity of the quantization distribution. A three-dimensional design space is explored, spanning bit-width allocation, group size configuration, and rotation strategies. Specifically, mixed-precision settings $\mathrm{K}_{2}\mathrm{V}_{2}$, $\mathrm{K}_{2}\mathrm{V}_{4}$, $\mathrm{K}_{4}\mathrm{V}_{2}$, and $\mathrm{K}_{4}\mathrm{V}_{4}$; key and value group sizes in $\{32,64,128\}$; and four rotation strategies (no rotation, key-only, value-only, and both). The evaluation encompasses generative tasks including CoQA, GSM8K, EQ-Bench, and LongBench, utilizing the same seven models as in quantization-error evaluation.
4.2 Mixed-Precision Quantization Error
Table 1 reports reconstruction errors at 2-, 3-, and 4-bit precision for seven representative models spanning multiple model families (Llama, Phi, Mistral, Qwen, DeepSeek) and datasets; full results appear in Appendix C.4. Figure 8 shows the complete error curves for Llama-3.3-70B on C4, and Figure 10 provides a per-layer breakdown for Llama-3.1-8B. Across all models, datasets, and bit-widths, key caches consistently incur larger reconstruction errors than value caches, and this gap remains stable across precision levels. These findings empirically support the theoretical prediction that key representations have higher energy and are therefore more sensitive to quantization.
4.3 Mixed-Precision Downstream Accuracy
Table 2 summarizes the Optimum Quanto results on GSM8K under both 1-shot and 8-shot Chain-of-Thought (CoT) prompting. Although CoT prompting can sometimes reduce reasoning accuracy, it is included here to assess the impact of longer contexts on quantized decoding. Across four representative models, including Llama-3.2-1B, Llama-3.1-8B, Phi-4-14B, and DeepSeek-R1-Qwen-14B, a consistent pattern emerges: allocating higher precision to the key cache ($\mathrm{K}_{4}\mathrm{V}_{2}$) preserves accuracy substantially better than the inverse ($\mathrm{K}_{2}\mathrm{V}_{4}$). On average, $\mathrm{K}_{4}\mathrm{V}_{2}$ recovers approximately 94% of the full-precision baseline, whereas value-favored allocations incur losses of up to 30 percentage points (pp). The performance gap widens with model scale; for instance, under 1-shot GSM8K, the $\mathrm{K}_{4}\mathrm{V}_{2}$ configuration outperforms $\mathrm{K}_{2}\mathrm{V}_{4}$ by 30 pp on Llama-3.2-1B and by 16 pp on Phi-4-14B. Notably, $\mathrm{K}_{4}\mathrm{V}_{2}$ nearly matches the symmetric $\mathrm{K}_{4}\mathrm{V}_{4}$ baseline despite halving the value bit budget, indicating that downstream performance is primarily constrained by key precision.
Table 2: GSM8K – Downstream accuracy with Optimum Quanto (token-wise) mixed-precision KV quantization. Token-wise quantization is applied with supported precisions $i,j\in\{2,4\}$, where $\mathrm{K}_{i}\mathrm{V}_{j}$ denotes $i$-bit keys and $j$-bit values. Results are shown for both 1-shot and 8-shot settings. Across models, the key-favored $\mathrm{K}_{4}\mathrm{V}_{2}$ consistently outperforms the value-favored $\mathrm{K}_{2}\mathrm{V}_{4}$ and approaches the $\mathrm{K}_{4}\mathrm{V}_{4}$ baseline, demonstrating the benefits of prioritizing key precision.
Model Shots $\mathbf{K}_{2}\mathbf{V}_{2}$ $\mathbf{K}_{2}\mathbf{V}_{4}$ $\mathbf{K}_{4}\mathbf{V}_{2}$ $\mathbf{K}_{4}\mathbf{V}_{4}$
Llama 3.2-1B-it 1-shot 0.033 0.035 0.338 0.357
8-shot 0.031 0.031 0.289 0.369
Llama 3.1-8B-it 1-shot 0.511 0.547 0.752 0.754
8-shot 0.408 0.441 0.770 0.782
Phi 4-14B 1-shot 0.759 0.783 0.913 0.923
8-shot 0.771 0.815 0.927 0.931
DeepSeek R1Q-14B 1-shot 0.772 0.775 0.865 0.867
8-shot 0.763 0.792 0.876 0.875
Table 3: Downstream accuracy with HQQ (channel-wise) mixed-precision KV quantization across GSM8K, EQ-Parseable, and CoQA. Each cell compares $\mathbf{K}_{i}\mathbf{V}_{x}$ and $\mathbf{K}_{x}\mathbf{V}_{i}$ for $i\in\{1,2,4,6,8\}$, where $x$ represents the mean accuracy averaged over the other cache’s bit-widths $B=\{1,2,4,6,8\}$. This corresponds to contrasting “allocate $i$ bits to keys (values averaged over BB)” versus “allocate $i$ bits to values (keys averaged over BB).” Across model scales from 0.6B to 32B and multiple tasks, key-favored allocations consistently yield higher accuracy, with more pronounced gains observed at lower precision and persistent benefits at higher precision, demonstrating that the key-first advantage generalizes beyond a single model, task, and quantization backend.